A magnetic field B= 0.6T is directed upward through a circular loop of diameter 7 cm and 500 turns. The loop is initially horizontal, so it is perpendicular to the magnetic field. It rotates through a horizontal axis so that the plane of the loop is at 74° with the horizontal axis within 1 second. What is the magnitude of the induced emf? 


1
Expert's answer
2022-06-01T12:49:05-0400
"{\cal E}=-\frac{\Delta \Phi}{\Delta t}"

"{\cal E}=-\frac{NBA(\cos\theta_2-\cos\theta_1)}{\Delta t}"

"{\cal E}=-\frac{500*0.6*3.14*0.035^2(\cos74^\circ-\cos0^\circ)}{1}=0.84\:\rm V"